### v1

# Energy decay of nonlocal viscoelastic equations with nonlinear damping and polynomial nonlinearity

Abstract.
This paper is concerned with the energy decay of a viscoelastic variable coefficient wave equation with nonlocality in time as well as nonlinear damping and polynomial nonlinear terms. Using the Lyapunov method, we establish a polynomial energy decay for the solution under relatively weak assumptions regarding the kernel of the nonlocal term. More specifically, we improve the decay rate of the energy by additionally imposing a certain convexity assumption on the kernel. Several examples are provided to confirm such improvements to faster polynomial or even exponential decays.
Key words and phrases:
Energy decay, viscoelastic equation, polynomial nonlinearity, frictional damping.
1School of Mathematics and Statistics & Hubei Key Laboratory of Engineering Modeling and Scientific Computing, Huazhong University of Science and Technology, Wuhan 430074, China.
2Department of Mathematics, Kyoto University, Kitashirakawa-Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan.
Corresponding author. E-mail: liu.yikan.8z@kyoto-u.ac.jp
1. Introduction
Let $\mathbb{R}_{+}:=(0,\infty)$ and $\Omega\subset\mathbb{R}^{n}$ ($n\geq 1$) be a bounded domain with a smooth boundary $\partial\Omega$. The main purpose of this manuscript is to investigate the energy decay rate of the following initial-boundary value problem for a nonlocal viscoelastic wave equation with variable coefficients, a nonlinear damping term and a polynomial nonlinear term:
{$&u_{tt}-\mathrm{div}(A(x)\nabla u)+\int_{0}^{t}f(t-s)\,\mathrm{div}(a(x)\nabla u(s))\,\mathrm{d}s\\ &\quad+b(x)h(u_{t})=k(x)|u|^{p-2}u\end{aligned}&\quad&\mbox{in }\Omega\times\mathbb{R}_{+},\\ &u=u^{0},\ u_{t}=u^{1}&\quad&\mbox{in }\Omega\times\{0\},\\ &u=0&\quad&\mbox{on }\partial\Omega\times\mathbb{R}_{+},\end{aligned}\right.$ (1.1)
where the initial displacement $u^{0}\in H_{0}^{1}(\Omega)$ and the initial velocity $u^{1}\in L^{2}(\Omega)$. The principal coefficient $A(x)$ is a symmetric and strictly positive-definite matrix-valued function on $\overline{\Omega}$. The nonlocal term $\int_{0}^{t}f(t-s)\,\mathrm{div}(a(x)\nabla u(s))\,\mathrm{d}s$ represents the memory effect in time, which takes the form of a convolution of the solution $u$ up to the second derivatives in space with the kernel function $f(t)$. Both of the nonlinear damping $h(u_{t})$ and polynomial nonlinear term $|u|^{p-2}u$ are equipped with $x$-dependent coefficients $b(x)$ and $k(x)$ respectively, among which we assume
$k\in C^{1}(\overline{\Omega}),\quad k\geq 0\mbox{ on }\overline{\Omega}.$
Moreover, the exponent constant $p$ satisfies
[…]
Detailed assumptions on $A(x),f(t),a(x),b(x)$ will be specified in Section 2.
Over the past few decades, significant results have been established regarding well-posedness and energy decays of nonlinear wave equations with either weak or strong damping. As a pioneering work, Webb [19] first investigated the well-posedness of solutions through the operator theory and examined the long-term behavior of these solutions using Lyapunov stability techniques for the semilinear wave equation with strong damping and a forcing term. Gazzola and Squassina [8] studied the semilinear wave equation with strong frictional damping and established both existence and non-existence results for solutions when the initial energy does not exceed the depth of the potential well. Further, they discovered the blow-up phenomena by potential well methods along with Levine’s concavity techniques. Subsequently, Ma and Fang [15] demonstrated both existence and non-existence of global weak solutions and derived energy decay estimates by combining potential well methods with a logarithmic Sobolev inequality applicable to the viscoelastic wave equations. Lian and Xu [13] further explored global existence, energy decay and infinite-time blow-up using methodologies similarly to [15]. Di, Shang and Song [6] discussed the semilinear wave equation with logarithmic nonlinearity and strong damping. They proved global well-posedness along with polynomial or exponential stability through potential methods combined with Lyapunov skills, and also considered various blow-up scenarios. Cui and Chai [5] investigated the energy decay for variable coefficient wave equations with logarithmic nonlinearity by applying vector assumptions and multiplicative methods. Moreover, Yang and Zhong [20] investigated the global well-posedness and blow-up phenomena for a strongly damped wave equation with a time-varying source and singular dissipation. They established both local well-posedness and globally existing solutions based on cut-off techniques, multiplier methods, contraction mapping principles and modified potential approaches. Meanwhile, they also derived the blow-up results of solutions with arbitrary positive initial energy and the lifespan of the blow-up solutions are derived.
Incorporating the nonlocal effect in time, viscoelastic wave equations with frictional damping have also witnessed significant developments concerning the energy decay and blow-up phenomena. In [4], Cavalcanti and Oquendo examined a viscoelastic equation with frictional damping as follows:
$u_{tt}-\triangle u+\int_{0}^{t}g(t-s)\,\mathrm{div}(a(x)\nabla u(s))\,\mathrm{d}s+b(x)h(u_{t})+f(u)=0.$
They derived decay rates when the kernel function $g(s)$ and the coefficients $a(x),b(x)$ satisfy certain conditions. Dias et al. [7] explored the energy decay of a similar wave equation in an open unbounded domain, where the coefficients $a(x),b(x)$ meet the aforementioned conditions and the kernel function satisfies some convexity conditions. Furthermore, Cavalcanti et al. [3] established stability results for an equation on a compact Riemannian manifold, with similar assumptions on $g,a,b$ similarly to those in [4]. Jin et al. [11] discussed uniform exponential and polynomial decay rates for solutions under weaker conditions than those in [4]. Mustafa and Abusharkh [17] investigated the energy decay of plate equations with frictional and viscoelastic damping, obtaining decay results under assumptions similarly to those in [7]. In particular, Li and Gao [12] used Lyapunov methods to derive blow-up results for viscoelastic equations and provided estimates for blow-up times. Ha and Park [9] also established blow-up results and local existence of solutions for viscoelastic equations with logarithmic nonlinearity using potential well methods and Lyapunov techniques. Additionally, Hao and Du [10] studied the energy decay and blow-up for viscoelastic equations with variable coefficients and logarithmic nonlinearity, employing potential well and Lyapunov methods akin to those in [9].
However, to the best of our knowledge, there are limited researches on a viscoelastic equation with variable coefficients that incorporates a frictional damping term $b(x)h(u_{t})$ and a nonlinear term $k(x)|u|^{p-2}u$. Motivated by the literature revisited above, in this article we investigate the energy decay rates of solutions to viscoelastic equations taking the form of (1.1). We establish energy decay results under weaker assumptions on the kernel function $f(t)$ compared to previous works such as [4, 3, 9, 10]. In this sense, our findings improve the aforementioned results.
The remainder of this paper is organized as follows. In Section 2, we specify assumptions on coefficients and prepare prelimineary lemmas including some elementary estimates of energies the global well-posedness of solutions. In Section 3, we state the two main results of this article, namely, the polynomial energy decay in Theorem 3.1 as well as a refined decay estimate in Theorem 3.2 under a weak convexity assumption on the kernel function $f(t)$. Further, we construct several examples to illustrate that Theorem 3.2 does imply faster decay rates in some cases. Subsequently, Sections 45 are devoted to the proofs of Theorems 3.13.2, respectively.
2. Preliminary and well-posedness
In this section, we fix notations and introduce some basic definitions, important lemmas and function spaces for the sake of the statements and proofs of our main results.
[…]
2.1. Some assumptions and definitions
We start with giving assumptions on the coefficients involved in the governing equation of (1.1).
(A1) The function $A(x)$ satisfies $A=(a_{ij})_{1\leq i,j\leq n}\in C^{\infty}(\overline{\Omega};\mathbb{R}^{n\times n})$ and there exist constants $\mu_{0}>\lambda_{0}>0$ such that
$\lambda_{0}|\xi|^{2}\leq A(x)\xi\cdot\,\xi\leq\mu_{0}|\xi|^{2},\quad\forall\,x\in\overline{\Omega},\ \forall\,\xi\in\mathbb{R}^{n}.$ (2.1)
(A2) The functions $a(x)$ and $b(x)$ satisfy
$0\leq a\in C^{1}(\overline{\Omega}),\quad 0\leq b\in L^{\infty}(\Omega),\quad a\not\equiv 0\mbox{ on }\partial\Omega$
and there exists a constant $\kappa>0$ such that
$a+b\geq\kappa>0\mbox{ in }\Omega.$
(A3) The function $h\in C(\mathbb{R})$ is non-decreasing and satisfies
$h(s)s\geq 0,\quad\forall\,s\in\mathbb{R}.$
Moreover, there exist constants $c_{i}>0$ ($i=1,2,3,4$) and $q\geq 1$ such that
$\begin{cases}c_{1}|s|\leq|h(s)|\leq c_{2}|s|,&|s|>1,\\ c_{3}|s|^{q}\leq|h(s)|\leq c_{4}|s|^{1/q},&|s|\leq 1.\end{cases}$
(A4) The function $f\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is non-increasing and satisfies
$f(0)>0,\quad\lambda_{0}-\|a\|_{\infty}\int_{0}^{\infty}f(t)\,\mathrm{d}t:=\ell>0.$ (2.2)
Next, we define
$F(t):=\int_{t}^{\infty}f(s)\,\mathrm{d}s,\quad M(\delta):=\int_{0}^{\infty}\frac{f(s)}{K_{\delta}(s)}\,\mathrm{d}s,\quad K_{\delta}(s):=-\frac{f^{\prime}(s)}{f(s)}+\delta,$ (2.3)
where $\delta\in(0,1)$ is a constant.
We introduce the energy functional as follows
$E(t)$ $:=\frac{1}{2}\|u_{t}(t)\|_{2}^{2}+\frac{1}{2}\int_{\Omega}A\nabla u(t)\cdot\nabla u(t)\,\mathrm{d}x-\frac{1}{2}\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x$
$\quad\>\,+\frac{1}{2}(f\circ\nabla u)(t)-\frac{1}{p}\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x,$ (2.4)
where
$(f\circ\nabla u)(t):=\int_{0}^{t}f(t-s)\int_{\Omega}a|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x\mathrm{d}s\geq 0.$ (2.5)
Here owing to assumptions (A1) and (A4), we estimate
$\quad\,\int_{\Omega}A\nabla u(t)\cdot\nabla u(t)\,\mathrm{d}x-\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x$
$\geq\int_{\Omega}\lambda_{0}|\nabla u(t)|^{2}\,\mathrm{d}x-\int_{0}^{\infty}f(t)\,\mathrm{d}t\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x\geq\left(\lambda_{0}-\|a\|_{\infty}\int_{0}^{\infty}f(t)\,\mathrm{d}t\right)\|\nabla u(t)\|_{2}^{2}$
$=\ell\|\nabla u(t)\|_{2}^{2}.$ (2.6)
To deal with the remaining negative term in (2.4), we define an auxiliary energy functional
$\mathbb{E}(t)$ $:=E(t)+\frac{1}{p}\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x$
$=\frac{1}{2}\|u_{t}(t)\|_{2}^{2}+\frac{1}{2}\int_{\Omega}A\nabla u(t)\cdot\nabla u(t)\,\mathrm{d}x-\frac{1}{2}\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x+\frac{1}{2}(f\circ\nabla u)(t).$ (2.7)
Then it is readily seen from (2.6) that
$\mathbb{E}(t)\geq\frac{\ell}{2}\|\nabla u(t)\|_{2}^{2}\geq 0.$ (2.8)
On the other hand, by formally differentiating (2.4) and employing the original problem (1.1), it is not difficult to calculate
$E^{\prime}(t)=\frac{1}{2}(f^{\prime}\circ\nabla u)(t)-\frac{f(t)}{2}\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x-\int_{\Omega}b\,h(u_{t}(t))u_{t}(t)\,\mathrm{d}x.$ (2.9)
Thanks to assumptions (A2)–(A4), we see that $E^{\prime}(t)\leq 0$, that is, the energy $E(t)$ is non-increasing.
2.2. Some lemmas and well-posedness
Lemma 2.1 (see [1]).
Let $r$ be a constant satisfying $2\leq r\leq\infty$ if $n=1,2$ and $2\leq r\leq 2_{*}=\frac{2n}{n-2}$ if $n\geq 3$. Then there is an optimal constant $B_{r}>0$ depending on $r$ such that
$\|w\|_{r}^{r}\leq B_{r}\|\nabla w\|_{2}^{r},\quad\forall\,w\in H_{0}^{1}(\Omega).$
Lemma 2.2.
Under assumptions (A1)–(A4), define
$H(\Lambda):=\frac{1}{2}\Lambda^{2}-\frac{KB}{p}\Lambda^{p},\quad\Lambda(t):=\left\{\ell\|\nabla u(t)\|_{2}^{2}+(f\circ\nabla u)(t)\right\}^{1/2},$
where $K:=\|k\|_{\infty},$ $B:=B_{p}\ell^{-p/2}$ and $\ell$ was the constant introduced in (2.2). Then there holds $E(t)\geq H(\Lambda(t))$.
Proof.
A direct application of Lemma 2.1 yields
$\frac{1}{p}\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x\leq\frac{K}{p}\|u(t)\|_{p}^{p}\leq\frac{KB_{p}}{p}\|\nabla u(t)\|_{2}^{p}.$ (2.10)
Then we estimate
$E(t)$ $\geq\frac{\ell}{2}\|\nabla u(t)\|_{2}^{2}+\frac{1}{2}(f\circ\nabla u)(t)-\frac{KB_{p}}{p}\|\nabla u(t)\|_{2}^{p}$
$\geq\frac{1}{2}\left\{\ell\|\nabla u(t)\|_{2}^{2}+(f\circ\nabla u)(t)\right\}-\frac{KB}{p}\left\{\ell\|\nabla u(t)\|_{2}^{2}+(f\circ\nabla u)(t)\right\}^{p/2}$
$=\frac{1}{2}\Lambda(t)^{2}-\frac{KB}{p}\Lambda(t)^{p}=H(\Lambda(t)).$
It is easy to verify that $H(\Lambda)$ attains a maximum at
$\Lambda_{1}:=\left(\frac{1}{KB}\right)^{\frac{1}{p-2}}$
with the maximum
$E_{1}:=H(\Lambda_{1})=\left(\frac{1}{2}-\frac{1}{p}\right)\Lambda_{1}^{2}>0.$
Lemma 2.3 (see [14, Lemma 3.4]).
Under assumptions (A1)–(A4), let $u$ be the solution to problem (1.1) with the initial data satisfying
$E(0)<E_{1},\quad\Lambda(0)<\Lambda_{1}.$
Then there exists a constant $\Lambda_{2}\in(0,\Lambda_{1})$ such that
$\Lambda(t)=\left\{\ell\|\nabla u(t)\|_{2}^{2}+(f\circ\nabla u)(t)\right\}^{1/2}\leq\Lambda_{2},\quad\forall\,t\in[0,T),$
where $T$ denotes the maximal existence time of the solution.
Next, we give the existence result of solutions to system (1.1).
Proposition 2.1 (Local existence).
Let $(u^{0},u^{1})\in H_{0}^{1}(\Omega)\times L^{2}(\Omega)$ be given and assumptions (A1)–(A4) hold. Then there exists $T>0$ such that problem (1.1) admits a unique local weak solution $u$ on $\overline{\Omega}\times[0,T]$ such that
$u\in C^{1}([0,T];L^{2}(\Omega))\cap C([0,T];H_{0}^{1}(\Omega)).$
By the Faedo-Galerkin method and the contraction mapping theorem, it is easy to prove Proposition 2.1. For detailed proof, please refer to [9] and [20].
Lemma 2.4.
[…]
$\frac{1}{p}\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x\leq\widetilde{C}E(t)\leq\widetilde{C}E(0),$ (2.11)
$\mathbb{E}(t)\leq(1+\widetilde{C})E(t)\leq(1+\widetilde{C})E(0),$ (2.12)
for any $t\in[0,T),$ where
$\widetilde{C}:=\frac{2KB\Lambda_{2}^{p-2}}{p-2KB\Lambda_{2}^{p-2}}>0.$
Proof.
First, by
$0<2KB\Lambda_{2}^{p-2}<2KB\Lambda_{1}^{p-2}=2<p,$
we confirm that the above $\widetilde{C}>0$. Then it directly follows from (2.10) that
$\frac{1}{p}\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x$ $\leq\frac{KB_{p}}{p}\|\nabla u(t)\|_{2}^{p}\leq\frac{KB}{p}\left(\ell^{1/2}\|\nabla u(t)\|_{2}\right)^{p-2}\ell\|\nabla u(t)\|_{2}^{2}$
$\leq\frac{2KB\Lambda_{2}^{p-2}}{p}\left(E(t)+\frac{1}{p}\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x\right),$
which implies the first half of (2.11). Combining (2.11) and (2.7), we obtain
$\mathbb{E}(t)=E(t)+\frac{1}{p}\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x\leq(1+\widetilde{C})E(t).$
The second halves of (2.11) and (2.12) are trivial due to the monotonicity of the energy $E(t)$. ∎
Remark 2.1.
Clearly, (2.12) indicates that $\mathbb{E}(t)$ is uniformly bounded for all $t\in[0,T)$. This implies that the solution is global and the maximal existence time $T=\infty$, for detailed proof please refer to [18]. At the same time, we have $0\leq E(t)\leq E(0)$ for all $t\in\overline{\mathbb{R}_{+}}$.
3. Main results and examples
In this section, we given the energy decay rates of global solutions of problem (1.1). The first result is as follows.
Theorem 3.1.
Suppose that assumptions (A1)–(A4) and the conditions in Lemma 2.3 hold. Let
$E(0)<\min\left\{E_{1},\left(\frac{\ell}{8KB_{p}}\right)^{\frac{2}{p-2}}\frac{\ell}{2(1+\widetilde{C})}\right\}.$ (3.1)
Then there exists a constant $C>0$ such that for any $t\geq t_{0}>0,$ the energy of problem (1.1) satisfies the polynomial decay
$\int_{0}^{\infty}E(t)^{\frac{q+1}{2}}\,\mathrm{d}t\leq CE(0),$ (3.2)
$E(t)\leq CE(0)(1+t)^{-\frac{2}{q+1}}.$ (3.3)
Below, we add an extra assumption on the kernel function $f(t)$ in addition to (A4).
(A5) Under assumption (A4), the function $f(t)$ further satisfies the following ordinary differential inequality
$f^{\prime}(t)\leq-\xi(t)G(f(t)),\quad t>0.$
Here $\xi\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is non-increasing and $G\in C^{2}([0,f(0)];\overline{\mathbb{R}_{+}})$ is a strictly increasing and strictly convex function satisfying $G(0)=G^{\prime}(0)=0$.
Remark 3.1 (see [16]).
If $G$ fulfills the above assumptions, then it allows an extension $\overline{G}$ in $\overline{\mathbb{R}_{+}}$ inheriting the same monotonicity, convexity and regularity. For example, if
$G(f(0))=g_{0},\quad G^{\prime}(f(0))=g_{1},\quad G^{\prime\prime}(f(0))=g_{2},$
then we can define $\overline{G}$ for $t>f(0)$ by
$\overline{G}(t)=\left(g_{0}-g_{1}f(0)+\frac{g_{2}}{2}f(0)^{2}\right)+(g_{1}-g_{2}f(0))t+\frac{g_{2}}{2}t^{2}.$
With slight abuse of the notation, we will still denote such an extension by $G$. Meanwhile, the strict monotonicity of $G$ guarantees the existence of the inverse function $G^{-1}:\overline{\mathbb{R}_{+}}\longrightarrow\overline{\mathbb{R}_{+}}$.
Theorem 3.2.
[…]
$\frac{1}{\varepsilon_{1}}G_{1}^{-1}\left(k_{1}\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s\right),$ $q=1,$ (3.4)
$E(0)\left\{t\,G_{2}^{-1}\left(\frac{k_{2}}{t\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s}\right)\right\}^{\frac{2}{q+1}},$ $q>1,$ (3.5)
for all $t>t_{0},$ where
$G_{1}(t):=\int_{t}^{f(0)}\frac{1}{s\,G^{\prime}(s)}\,\mathrm{d}s,\quad G_{2}(t):=t\,G^{\prime}(\varepsilon_{1}t).$
Remark 3.2.
According to the assumptions on $G$ in (A5) and the above definitions of $G_{1},G_{2}$, we see that $G_{1}$ is strictly decreasing and defines a bijection from $[0,f(0)]$ to $\overline{\mathbb{R}_{+}}$, and $G_{2}$ is strictly increasing and defines a bijection from $\overline{\mathbb{R}_{+}}$ to $\overline{\mathbb{R}_{+}}$.
Remark 3.3.
Theorems 3.13.2 still hold for $p=2$, i.e., the polynomial nonlinearity in (1.1) degenerates to the linear one. The proofs are rather analogous and thus we omit them here.
Next, we use three examples to illustrate Theorem 3.2.
Example 3.1.
Let the kernel function $f(t)$ takes the form of $f(t)=\alpha\,e^{-\beta(1+t)}$, where $\alpha,\beta>0$ are sufficiently small constants so that $f$ satisfies (A4). Then $f$ satisfies (A5) with $\xi(t)=\beta$ and $G(t)=t$. Therefore, it follows from Theorem 3.2 and direct calculations that
$E(t)\leq\begin{cases}c_{1}\mathrm{e}^{-c_{2}t},&q=1,\\ c_{3}t^{-\frac{2}{q-1}},&q>1,\end{cases}$
where $c_{1},c_{2},c_{3}>0$ are some constants.
Example 3.2.
Let $q=1$ and $f(t)=\alpha\exp(-t^{\beta})$, where $\beta\in(0,1)$ and $a>0$ are again small constants so that $f$ satisfies (A4). Then this time $f$ satisfies (A5) with $\xi(t)=1$ and
$G(t)=\frac{\beta\,t}{(\ln(\alpha/t))^{1/\beta-1}}.$
By
$G^{\prime}(t)=\frac{(1-\beta)+\beta\ln(\alpha/t)}{(\ln(\alpha/t))^{1/\beta}},\quad G^{\prime\prime}(t)=\frac{(1-\beta)(\ln(\alpha/t)+1/\beta)}{(\ln(\alpha/t))^{\frac{1}{\beta+1}}},$
we see that $G$ also achieves the requirement in (A5) on $[0,\alpha]$. Then we deal with $G_{1}$ as
$G_{1}(t)$ $=\int_{t}^{\alpha}\frac{1}{s\,G^{\prime}(s)}\,\mathrm{d}s=\int_{t}^{\alpha}\frac{(\ln(\alpha/s))^{1/\beta}}{s(1-\beta+\beta\ln(\alpha/s))}\,\mathrm{d}s=\int_{0}^{\ln(\alpha/t)}\frac{\tau^{1/\beta}}{1-\beta+\beta\tau}\,\mathrm{d}\tau$
$=\frac{1}{\beta}\int_{0}^{\ln(\alpha/t)}{\tau^{1/\beta-1}}\left(\frac{\tau}{\beta^{-1}-1+\tau}\right)\mathrm{d}\tau\leq\frac{1}{\beta}\int_{0}^{\ln(\alpha/t)}\tau^{1/\beta-1}\,\mathrm{d}\tau=\left(\ln\frac{\alpha}{t}\right)^{1/\beta}.$
Then it follows from (3.4) that $E(t)\leq c\exp(-c\,t^{\beta})$ with some constant $c>0$.
Example 3.3.
Now let us consider $q>1$ and $f(t)=\alpha(1+t)^{-\beta}$, where $\alpha>0$ and $\beta>1$ are suitably chosen so that (A4) holds. Similarly as before, $f$ satisfies (A5) with $\xi(t)=1$ and $G(t)=t^{\frac{\beta+1}{\beta}}$. Then again $G$ achieves the condition in (A5) hold and it follows from (3.5) that
$E(t)\leq c\,t^{\min\{-\frac{2}{q+1},\frac{2(q-1)}{(q+1)^{2}}\}},$
with some constant $c>0$.
From the above examples, it is evident that the rates of energy decay can be significantly faster than those stated in Theorem 3.1 when $q=1$ and $G(t)$ is linear in the case of $q>1$. However, when $G(t)$ is nonlinear in the case of $q>1$, the rate of energy decay of the solution does not appear to increase substantially due to the weak dissipation conditions.
4. Proof of Theorem 3.1
In the sequel, by $C>0$ we denote generic constants which may change from line to line. In order to prove Theorem 3.1, we need to introduce some auxiliary functions and lemmas. First we define
$F_{1}(t):=\int_{0}^{t}F(t-s)\int_{\Omega}a|\nabla u(s)|^{2}\,\mathrm{d}x\mathrm{d}s,\quad F_{2}(t):=M(\delta)(\delta F_{1}(t)+E(t)),$
where $F(t),M(\delta)$ were defined in (2.3). We recall the following result.
Lemma 4.1 (see [11, Lemma 2.1]).
For $t\geq 0,$ the functions $F_{1}(t),F_{2}(t)$ defined above satisfy
$\frac{\mathrm{d}}{\mathrm{d}t}F_{1}(t)$ $\leq-\frac{1}{2}(f\circ\nabla u)(t)+2\lambda_{0}\|\nabla u(t)\|_{2}^{2},$ (4.1)
$\frac{\mathrm{d}}{\mathrm{d}t}F_{2}(t)$ $\leq-\frac{1}{2}M(\delta)\int_{0}^{t}K_{\delta}(t-s)f(t-s)\int_{\Omega}a|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x\mathrm{d}s+2\lambda_{0}M(\delta)\delta\|\nabla u(t)\|_{2}^{2},$ (4.2)
where $(f\circ\nabla u)(t)$ was defined in (2.5) and $\lambda_{0}$ was the constant in (2.1). Moreover, we have
$M(\delta)\delta\longrightarrow 0\quad\mbox{as }\delta\to 0.$ (4.3)
Next, let us construct a nonnegative function $\widetilde{a}\in C^{1}(\overline{\Omega})$ satisfying $\mathrm{supp}\,\widetilde{a}\subset\mathrm{supp}\,a$ and
$\widetilde{a}(x)=\begin{cases}a(x)&\mbox{if }a(x)\geq\kappa/2,\\ 0&\mbox{if }a(x)\leq\kappa/4.\end{cases}$
Then assumption (A2) gives
$\widetilde{a}+b\geq\frac{\kappa}{2}\quad\mbox{in }\Omega.$ (4.4)
We invoke a useful result.
Lemma 4.2 (see [1]).
Let $\Omega_{1},\Omega_{2}$ be subsets of $\Omega$ with positive measures such that
$\overline{\Omega}_{1}\subset\Omega_{2}\subset\Omega,\quad\mathrm{meas}(\partial\Omega_{2}\cap\partial\Omega)>0.$
Then there exists a constant $C_{0}>0$ such that
$\int_{\Omega_{1}}|w|^{2}\,\mathrm{d}x\leq C_{0}\int_{\Omega_{2}}|\nabla w|^{2}\,\mathrm{d}x,\quad\forall\,w\in H_{0}^{1}(\Omega).$
According to assumption (A2), there exist a constant $\varepsilon_{0}>0$ and a subset $V\subset\overline{\Omega}$ such that
$\mathrm{meas}(\partial V\cap\partial\Omega)>0,\quad a\geq\varepsilon_{0}\mbox{in }V.$
Setting
$\Omega_{1}:=\mathrm{supp}\,\widetilde{a},\quad\Omega_{2}:=\{x\in\Omega\mid a(x)>\min\{\kappa/4,\varepsilon_{0}\}\},$
we see that $\Omega_{1},\Omega_{2}$ satisfy the assumption of Lemma 4.2. Then for any $w\in H_{0}^{1}(\Omega)$, we employ Lemma 4.2 to deduce
$\int_{\Omega}\widetilde{a}|w|^{2}\,\mathrm{d}x$ $=\int_{\Omega_{1}}\widetilde{a}|w|^{2}\,\mathrm{d}x\leq\|\widetilde{a}\|_{\infty}\int_{\Omega_{1}}|w|^{2}\,\mathrm{d}x\leq\|\widetilde{a}\|_{\infty}C_{0}\int_{\Omega_{2}}|\nabla w|^{2}\,\mathrm{d}x$
$\leq C\int_{\Omega_{2}}a|\nabla w|^{2}\,\mathrm{d}x\leq C\int_{\Omega}a|\nabla w|^{2}\,\mathrm{d}x.$ (4.5)
Now we introduce further auxiliary functions
$F_{3}(t):=\int_{\Omega}u_{t}(t)u(t)\,\mathrm{d}x,\quad F_{4}(t):=-\int_{0}^{t}f(t-s)\int_{\Omega}\widetilde{a}\,u_{t}(t)(u(t)-u(s))\,\mathrm{d}x\mathrm{d}s.$
In the next two lemmas, we provide estimates for the derivatives of the above functions.
Lemma 4.3.
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}F_{3}(t)$ $\leq\|u_{t}(t)\|_{2}^{2}-\frac{\ell}{2}\|\nabla u(t)\|_{2}^{2}+\frac{\|a\|_{\infty}}{\ell}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(s)-u(t))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\quad\,+\frac{\|b\|_{\infty}B_{2}}{\ell}\int_{\Omega}b\,h(u_{t}(t))^{2}\,\mathrm{d}x+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x,$ (4.6)
where $\ell$ and $B_{2}$ were the constants introduced in (2.2) and Lemma 2.1 with $r=2,$ respectively.
Proof.
By the governing equation and the homogeneous boundary condition in (1.1), we utilize the divergence theorem to calculate
$\frac{\mathrm{d}}{\mathrm{d}t}F_{3}(t)$ $=\int_{\Omega}\left\{u_{tt}(t)u(t)+u_{t}(t)^{2}\right\}\mathrm{d}x$
$=\|u_{t}(t)\|_{2}^{2}+\int_{\Omega}\left\{\mathrm{div}(A\nabla u(t))-\int_{0}^{t}f(t-s)\,\mathrm{div}(a\nabla u(s))\,\mathrm{d}s-b\,h(u_{t}(t))\right\}u(t)\,\mathrm{d}x$
$\quad\,+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x$
$=\|u_{t}(t)\|_{2}^{2}+I_{0}(t)+I_{1}(t)+I_{2}(t)+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x,$ (4.7)
where
$I_{0}(t):=-\int_{\Omega}A\nabla u(t)\cdot\nabla u(t)\,\mathrm{d}x\leq-\lambda_{0}\|\nabla u(t)\|_{2}^{2},$ (4.8)
by assumption (A1), and
$I_{1}(t):=\int_{\Omega}a\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla u(s)\,\mathrm{d}s\mathrm{d}x,\quad I_{2}(t):=-\int_{\Omega}b\,h(u_{t}(t))u(t)\,\mathrm{d}x.$
For $I_{1}(t)$, we separate $u(s)=u(t)+(u(s)-u(t))$ and use assumptions (A2) and (A4) to estimate
$I_{1}(t)$ $=\int_{\Omega}a\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla u(t)\,\mathrm{d}s\mathrm{d}x+\int_{\Omega}a\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla(u(s)-u(t))\,\mathrm{d}s\mathrm{d}x$
$=\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x+\int_{\Omega}\sqrt{a}\,\nabla u(t)\cdot\sqrt{a}\int_{0}^{t}f(t-s)\nabla(u(s)-u(t))\,\mathrm{d}s\mathrm{d}x$ (4.9)
$\leq\int_{0}^{\infty}f(t)\,\mathrm{d}t\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x+\epsilon_{1}\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x$
$\quad\,+\frac{1}{4\epsilon_{1}}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(s)-u(t))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\|a\|_{\infty}\left(\int_{0}^{\infty}f(t)\,\mathrm{d}t+\epsilon_{1}\right)\|\nabla u(t)\|_{2}^{2}+\frac{1}{4\epsilon_{1}}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(s)-u(t))\,\mathrm{d}s\right|^{2}\mathrm{d}x.$ (4.10)
Here we applied Cauchy’s inequality to treat the second term in (4.9), where $\epsilon_{1}>0$ is a parameter to be determined later. In the same manner, we estimate $I_{2}(t)$ as
$I_{2}(t)$ $\leq\int_{\Omega}\sqrt{b}\,|h(u_{t}(t))|\sqrt{b}\,|u(t)|\,\mathrm{d}x\leq\epsilon_{2}\int_{\Omega}b\,u(t)^{2}\,\mathrm{d}x+\frac{1}{4\epsilon_{2}}\int_{\Omega}b\,h(u_{t}(t))^{2}\,\mathrm{d}x$
$\leq\epsilon_{2}\|b\|_{\infty}\|u(t)\|_{2}^{2}+\frac{1}{4\epsilon_{2}}\int_{\Omega}b\,h(u_{t}(t))^{2}\,\mathrm{d}x$
$\leq\epsilon_{2}\|b\|_{\infty}B_{2}\|\nabla u(t)\|_{2}^{2}+\frac{1}{4\epsilon_{2}}\int_{\Omega}b\,h(u_{t}(t))^{2}\,\mathrm{d}x$ (4.11)
again with a parameter $\epsilon_{2}>0$ to be determined later, where we utilized Lemma 2.1 with $r=2$ to deduce the last inequality.
Recalling the definition of $\ell$ in (2.2), we substitute inequalities (4.8), (4.10) and (4.11) into (4.7) to derive
$\frac{\mathrm{d}}{\mathrm{d}t}F_{3}(t)$ $\leq\|u_{t}(t)\|_{2}^{2}-\left|\ell-\|a\|_{\infty}\epsilon_{1}-\|b\|_{\infty}B_{2}\epsilon_{2}\right|\|\nabla u(t)\|_{2}^{2}+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x$
$\quad\,+\frac{1}{4\epsilon_{1}}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(s)-u(t))\,\mathrm{d}s\right|^{2}\mathrm{d}x+\frac{1}{4\epsilon_{2}}\int_{\Omega}b\,h(u_{t}(t))^{2}\,\mathrm{d}x$
Consequently, we arrive at (4.6) by simply choosing
$\epsilon_{1}=\frac{\ell}{4\|a\|_{\infty}},\quad\epsilon_{2}=\frac{\ell}{4\|b\|_{\infty}B_{2}},$
which completes the proof. ∎
Lemma 4.4.
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}F_{4}(t)$ $\leq\varepsilon\|u_{t}(t)\|_{2}^{2}+\varepsilon(1+B_{2(p-1)}\eta_{1})\|\nabla u(t)\|_{2}^{2}+C(\varepsilon)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\quad\,-\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}\widetilde{a}\,u_{t}(t)^{2}\,\mathrm{d}x+C(\varepsilon)\int_{\Omega}b\,h^{2}(u_{t}(t))\,\mathrm{d}x-\frac{f(0)C}{4\varepsilon}(f^{\prime}\circ\nabla u)(t).$ (4.12)
Here $\varepsilon\in(0,1)$ is an arbitrary constant, $C(\varepsilon)>0$ is a constant depending on $\varepsilon$ and
$\eta_{1}:=\left(\frac{2(1+\widetilde{C})E(0)}{\ell}\right)^{p-2},$
where $\widetilde{C}$ was the constant introduced in Lemma 2.4.
Proof.
Similarly to the proof of Lemma 4.3, we differentiate $F_{4}(t)$ by its definition and exploit problem (1.1) to calculate
$\frac{\mathrm{d}}{\mathrm{d}t}F_{4}(t)$ $=-\int_{\Omega}u_{tt}(t)\widetilde{a}\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x$
$\quad\,-\int_{\Omega}\widetilde{a}\,u_{t}(t)\int_{0}^{t}f^{\prime}(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x-\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}\widetilde{a}\,u_{t}(t)^{2}\,\mathrm{d}x$
$=\int_{\Omega}\left\{-\mathrm{div}(A\nabla u(t))+\int_{0}^{t}f(t-s)\,\mathrm{div}(a\nabla u(s))\,\mathrm{d}s+b\,h(u_{t}(t))-k\,|u(t)|^{p-2}u(t)\right\}$
×$\qquad\>\>\times\widetilde{a}\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x+I_{9}(t)+I_{10}(t)$
$=\sum_{i=3}^{10}I_{i}(t),$
where
$I_{3}(t)$ $:=\int_{\Omega}\widetilde{a}\,A\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x,$
$I_{4}(t)$ $:=\int_{\Omega}\nabla\widetilde{a}\cdot A\nabla u(t)\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x,$
$I_{5}(t)$ $:=-\int_{\Omega}\widetilde{a}\int_{0}^{t}f(t-s)a\nabla u(s)\,\mathrm{d}s\cdot\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x,$
$I_{6}(t)$ $:=-\int_{\Omega}\nabla\widetilde{a}\cdot\int_{0}^{t}f(t-s)a\nabla u(s)\,\mathrm{d}s\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x,$
$I_{7}(t)$ $:=\int_{\Omega}b\,h(u_{t}(t))\left(\int_{0}^{t}f(t-s)\widetilde{a}(u(t)-u(s))\,\mathrm{d}s\right)\mathrm{d}x,$
$I_{8}(t)$ $:=-\int_{\Omega}\widetilde{a}\,k|u(t)|^{p-2}u(t)\left(\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\right)\mathrm{d}x,$
$I_{9}(t)$ $:=-\int_{\Omega}\widetilde{a}\,u_{t}(t)\int_{0}^{t}f^{\prime}(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x,\quad I_{10}(t):=-\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}\widetilde{a}\,u_{t}(t)^{2}\,\mathrm{d}x.$
We notice that $I_{10}(t)$ already appears in the desired estimate (4.12). Then we shall estimate the remaining terms above one by one.
Using Young’s inequality and (4.5), one has
$|I_{3}(t)|$ $\leq\int_{\Omega}\left|\sqrt{\widetilde{a}}\,\nabla u(t)\cdot\sqrt{\widetilde{a}}\,A\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|\mathrm{d}x$
$\leq\frac{\varepsilon}{4\|\widetilde{a}\|_{\infty}}\int_{\Omega}\widetilde{a}\,|\nabla u(t)|^{2}\,\mathrm{d}x+\frac{\|\widetilde{a}\|_{\infty}}{\varepsilon}\int_{\Omega}\widetilde{a}\left|A\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+\frac{\|\widetilde{a}\|_{\infty}\widetilde{\lambda}}{\varepsilon}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x,$ (4.13)
where $\widetilde{\lambda}:=\max_{x\in\overline{\Omega}}\|A(x)\|_{\ell^{2}}^{2}$ and $\|A(x)\|_{\ell^{2}}$ denotes the matrix $2$-norm of $A(x)\in\mathbb{R}^{n\times n}$, i.e., $|A(x)\xi|\leq\|A(x)\|_{\ell^{2}}|\xi|$ for any $\xi\in\mathbb{R}^{n}$. In the same manner and employing the definition of $\widetilde{a}$ and (4.5), we deduce that
$|I_{4}(t)|$ $\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+\frac{1}{\varepsilon}\int_{\Omega}\left|\nabla\widetilde{a}\,A\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+\frac{\|\nabla a\|_{\infty}^{2}\widetilde{\lambda}}{\varepsilon}\int_{\Omega_{1}}\left|\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+\frac{C\|\nabla a\|_{\infty}^{2}\widetilde{\lambda}}{\varepsilon}\int_{\Omega_{2}}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+\frac{C\|\nabla a\|_{\infty}^{2}\widetilde{\lambda}}{\varepsilon}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x.$
For $I_{5}(t)$, we separate $u(s)=u(t)+(u(s)-u(t))$ and use assumptions (A2) and (A4) to estimate
$|I_{5}(t)|$ $=\left|\int_{\Omega}\widetilde{a}\int_{0}^{t}f(t-s)a\nabla(u(s)-u(t)+u(t))\,\mathrm{d}s\cdot\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x\right|$
$\leq\|\widetilde{a}\|_{\infty}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\quad\,+\|\widetilde{a}\|_{\infty}\int_{0}^{t}f(s)\,\mathrm{d}s\int_{\Omega}a\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x$
$\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+\|\widetilde{a}\|_{\infty}\left(1+\frac{(\lambda_{0}-\ell)^{2}}{\varepsilon\|a\|_{\infty}}\right)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+C(\varepsilon)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x.$ (4.14)
Similarly, we estimate $I_{6}(t)$ as
$|I_{6}(t)|\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+C(\varepsilon)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x.$ (4.15)
In the same manner, we estimate $I_{7}(t)$ and $I_{9}(t)$ as
$|I_{7}(t)|$ $\leq\varepsilon\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x+C(\varepsilon)\int_{\Omega}b\,h(u_{t}(t))^{2}\,\mathrm{d}x,$ (4.16)
$|I_{9}(t)|$ $\leq\varepsilon\|u_{t}(t)\|_{2}^{2}+\frac{1}{4\varepsilon}\int_{\Omega}\left|\widetilde{a}\int_{0}^{t}f^{\prime}(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\varepsilon\|u_{t}(t)\|_{2}^{2}+\frac{C}{4\varepsilon}\int_{\Omega}a\int_{0}^{t}f^{\prime}(t-s)\,\mathrm{d}s\int_{0}^{t}f^{\prime}(t-s)|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}s\mathrm{d}x$
$\leq\varepsilon\|u_{t}(t)\|_{2}^{2}-\frac{f(0)C}{4\varepsilon}(f^{\prime}\circ\nabla u)(t).$ (4.17)
Finally, for $I_{8}(t)$, we utilize Young’s inequality and (4.5) to derive
$|I_{8}(t)|$ $\leq\varepsilon\int_{\Omega}|u(t)|^{2(p-1)}\,\mathrm{d}x+\frac{1}{4\varepsilon}\int_{\Omega}k^{2}\,\widetilde{a}^{2}\left|\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\varepsilon\|u(t)\|_{2(p-1)}^{2(p-1)}+\frac{K^{2}C\|\widetilde{a}\|_{\infty}}{4\varepsilon}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\varepsilon B_{2(p-1)}\|\nabla u(t)\|_{2}^{2(p-1)}+C(\varepsilon)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\varepsilon B_{2(p-1)}\eta_{1}\|\nabla u(t)\|_{2}^{2}+C(\varepsilon)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x,$ (4.18)
where we used Lemma 2.1 with $r=2(p-1)$ and recall $K=\|k\|_{\infty}$. For the first term in (4.18), we combine (2.12) with (2.8) to deduce
$\|\nabla u(t)\|_{2}^{2(p-2)}\leq\left(\frac{2\,\mathbb{E}(t)}{\ell}\right)^{p-2}\leq\left(\frac{2(1+\widetilde{C})E(0)}{\ell}\right)^{p-2}=\eta_{1}.$
Eventually, we arrive at (4.12) by combining all above estimates (4.13)–(4.18). ∎
We are now prepared to demonstrate Theorem 3.1. First we define an auxiliary function
$L(t):=N_{1}E(t)+F_{3}(t)+N_{2}F_{4}(t),$
where $N_{1},N_{2}>0$ are constants to be determined soon. Using Young’s inequality and (4.5), we easily obtain
$|F_{3}(t)|$ $\leq\frac{B_{2}}{2}\|\nabla u(t)\|_{2}^{2}+\frac{1}{2}\|u_{t}(t)\|_{2}^{2},$
$|F_{4}(t)|$ $\leq\frac{1}{2}\|u_{t}(t)\|_{2}^{2}+\frac{\|\widetilde{a}\|_{\infty}}{2}\int_{\Omega}\widetilde{a}\left|\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\frac{1}{2}\|u_{t}(t)\|_{2}^{2}+\frac{C\|\widetilde{a}\|_{\infty}}{2}\int_{\Omega}a\int_{0}^{t}f(t-s)\,\mathrm{d}s\int_{0}^{t}f(t-s)|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}s\mathrm{d}x$
$\leq\frac{1}{2}\|u_{t}(t)\|_{2}^{2}+\frac{C\lambda_{0}}{2}\int_{\Omega}a\int_{0}^{t}f(t-s)|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}s\mathrm{d}x$
$=\frac{1}{2}\|u_{t}(t)\|_{2}^{2}+\frac{C\lambda_{0}}{2}(f\circ\nabla u)(t).$
Then by choosing $N_{1}$ and $N_{2}$ suitably, we have the equivalence $L(t)\sim E(t)$.
Fix any $t_{0}>0$ and put $f_{1}:=\int_{0}^{t_{0}}f(s)\,\mathrm{d}s$. Combining (2.9), Lemma 4.3 and Lemma 4.4 with
$\varepsilon=\frac{\ell}{4N_{2}(1+B_{2(p-1)}\eta_{1})},$
we obtain for all $t\geq t_{0}$ that
$L^{\prime}(t)$ $\leq-\frac{\ell}{4}\|\nabla u(t)\|_{2}^{2}+(\varepsilon N_{2}+1)\|u_{t}(t)\|_{2}^{2}-N_{2}f_{1}\int_{\Omega}(\widetilde{a}+b)u_{t}(t)^{2}\,\mathrm{d}x+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x$
$\quad\,+(C(\varepsilon)N_{2}+C)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s)\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\quad\,+\left(\frac{N_{1}}{2}-\frac{f(0)CN_{2}}{4\varepsilon}\right)(f^{\prime}\circ\nabla u)(t)+(C(\varepsilon)N_{2}+C)\int_{\Omega}b(h(u_{t}(t))^{2}+u_{t}(t)^{2})\,\mathrm{d}x.$
At this point, we first use (4.4) and choose $N_{2}$ sufficiently large so that
$\theta:=N_{2}f_{1}\frac{\kappa}{2}-\frac{\ell}{4(1+B_{2(p-1)}\eta_{1})}-1>0,$
and then choose $N_{1}$ sufficiently large so that
$\frac{N_{1}}{2}-\frac{f(0)CN_{2}}{4\varepsilon}>0.$
Hence we arrive at
$L^{\prime}(t)$ $\leq-\frac{\ell}{4}\|\nabla u(t)\|_{2}^{2}-\theta\int_{\Omega}u_{t}(t)^{2}\,\mathrm{d}x+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x+(C(\varepsilon)N_{2}+C)\int_{\Omega}b(h(u_{t}(t))^{2}+u_{t}(t)^{2})\,\mathrm{d}x$
$\quad\,+(C(\varepsilon)N_{2}+C)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s)\,\mathrm{d}s\right|^{2}\mathrm{d}x.$ (4.19)
For the last term above, we recall the definition (2.3) of $K_{\delta}(t)$ and $M(\delta)$ to estimate
$\quad\,\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s)\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$=\int_{\Omega}a\left|\int_{0}^{t}\sqrt{\frac{f(t-s)}{K_{\delta}(t-s)}}\sqrt{f(t-s)K_{\delta}(t-s)}\nabla(u(t)-u(s)\,\mathrm{d}s\right|^{2}\mathrm{d}x$
$\leq\int_{0}^{t}\frac{f(s)}{K_{\delta}(s)}\,\mathrm{d}s\int_{\Omega}a\int_{0}^{t}K_{\delta}(t-s)f(t-s)|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}s\mathrm{d}x$
$\leq M(\delta)\int_{0}^{t}K_{\delta}(t-s)f(t-s)\int_{\Omega}a|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x\mathrm{d}s.$ (4.20)
Next, we further define
$J(t):=L(t)+\frac{\ell}{32\lambda_{0}}F_{1}(t)+{2(C(\varepsilon)N_{2}+C)}F_{2}(t).$
Then we apply inequalities (4.1)–(4.2) in Lemma 4.1 and (4.20) to (4.19) to derive
$J^{\prime}(t)$ $\leq-\left(\frac{\ell}{4}-\frac{\ell}{16}-{4(C(\varepsilon)N_{2}+C)\lambda_{0}\delta M(\delta)}\right)\|\nabla u(t)\|_{2}^{2}-\theta\|u_{t}(t)\|_{2}^{2}-\frac{\ell}{64\lambda_{0}}(f\circ\nabla u)(t)$
$\quad\,+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x+(C(\varepsilon)N_{2}+C)\int_{\Omega}b(h(u_{t}(t))^{2}+u_{t}(t)^{2})\,\mathrm{d}x.$
Here, the convergence (4.3) guarantees the existence of a constant $\delta_{0}>0$ such that
$4(C(\varepsilon)N_{2}+C)\lambda_{0}\delta M(\delta)<\frac{\ell}{16},$
for any $0<\delta<\delta_{0}$. Recalling the original energy $E(t)$, we deduce
$J^{\prime}(t)$ $\leq-\frac{\ell}{8}\|\nabla u(t)\|_{2}^{2}-\theta\|u_{t}(t)\|_{2}^{2}-\frac{\ell}{64\lambda_{0}}(f\circ\nabla u)(t)+\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x$
$\quad\,+(C(\varepsilon)N_{2}+C)\int_{\Omega}b(h(u_{t}(t))^{2}+u_{t}(t)^{2})\,\mathrm{d}x$
$\leq-\epsilon\,E(t)-\left(\frac{\ell}{8}-\frac{\epsilon\mu_{0}}{2}\right)\|\nabla u(t)\|_{2}^{2}-\left(\theta-\frac{\epsilon}{2}\right)\|u_{t}(t)\|_{2}^{2}-\left(\frac{\ell}{64\lambda_{0}}-\frac{\epsilon}{2}\right)(f\circ\nabla u)(t)$
$\quad\,+\left(1-\frac{\epsilon}{p}\right)\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x+(C(\varepsilon)N_{2}+C)\int_{\Omega}b(h(u_{t}(t))^{2}+u_{t}(t)^{2})\,\mathrm{d}x,$
where $\epsilon>0$ is another adjustable parameter. Utilizing (2.8), (2.10) and (2.12), we bound
$\int_{\Omega}k|u(t)|^{p}\,\mathrm{d}x$ $\leq KB_{p}\|\nabla u(t)\|_{2}^{p}\leq KB_{p}\left(\frac{2\,\mathbb{E}(t)}{\ell}\right)^{\frac{p-2}{2}}\|\nabla u(t)\|_{2}^{2}$
$\leq KB_{p}\left(\frac{2(1+\widetilde{C})E(0)}{\ell}\right)^{\frac{p-2}{2}}\|\nabla u(t)\|_{2}^{2}.$ (4.21)
Thanks to the key assumption (3.1), it is possible to choose a sufficiently small $\epsilon>0$ such that
$\frac{\ell}{8}-\frac{\epsilon\mu_{0}}{2}-KB_{p}\left(\frac{2(1+\widetilde{C})E(0)}{\ell}\right)^{\frac{p-2}{2}}>0,\quad\theta-\frac{\epsilon}{2}>0,\quad\frac{\ell}{64\lambda_{0}}-\frac{\epsilon}{2}>0,$
and hence
$J^{\prime}(t)+\epsilon\,E(t)\leq C\int_{\Omega}b(h(u_{t}(t))^{2}+u_{t}(t)^{2})\,\mathrm{d}x.$ (4.22)
In the sequel, we discuss the cases of $q=1$ and $q>1$ separately.
Case 1. In the case of $q=1$, the assumption (A3) asserts that $h$ behaves linearly in $\mathbb{R}$. Then it follows from (4.22) that
$J^{\prime}(t)+\epsilon\,E(t)\leq C\int_{\Omega}b\,h(u_{t}(t))u_{t}(t)\,\mathrm{d}x\leq-CE^{\prime}(t),$ (4.23)
where we employed (2.9) to obtain the second inequality above. Picking any $\tau>t_{0}>0$, we integrate the above inequality over $(t_{0},\tau)$ to estimate
$J(\tau)-J(t_{0})+\epsilon\int_{t_{0}}^{\tau}E(t)\,\mathrm{d}t\leq-C\int_{t_{0}}^{\tau}E^{\prime}(t)\,\mathrm{d}t=C(E(t_{0})-E(\tau))\leq CE(0),$
due to the monotonicity of $E(t)$. Since $J(\tau)\geq 0$ and $J(t_{0})\leq CE(0)$ for some $t_{0}>0$, we conclude
$\int_{t_{0}}^{\tau}E(t)\,\mathrm{d}t\leq CJ(t_{0})+CE(0)\leq CE(0).$
Since $\tau>t_{0}$ was chosen arbitrarily, the above estimate holds also as $\tau\to\infty$. Hence we arrive at (3.2) by
$\int_{0}^{\infty}E(t)\,\mathrm{d}t=\left(\int_{0}^{t_{0}}+\int_{t_{0}}^{\infty}\right)E(t)\,\mathrm{d}t\leq t_{0}E(0)+CE(0)\leq CE(0),$
again by the monotonicity of $E(t)$, which further indicates (3.3) immediately.
Case 2. In the case of $q>1$, we apply [11, Equation (4.11)] to the right-hand side of (4.22) to derive
$J^{\prime}(t)+\epsilon\,E(t)\leq C\left\{-E^{\prime}(t)+(-E^{\prime}(t))^{\frac{2}{q+1}}\right\}.$ (4.24)
Now introduce an auxiliary function $\varphi:\overline{\mathbb{R}_{+}}\longrightarrow\overline{\mathbb{R}_{+}}$ which is non-increasing and locally absolutely continuous. For the constant $t_{0}>0$ in Case 1 and any $\tau>t_{0}$, we multiply (4.24) by $\varphi(t)$ and integrate over $(t_{0},\tau)$ to deduce
$\int_{t_{0}}^{\tau}\varphi(t)E(t)\,\mathrm{d}t\leq-C\int_{t_{0}}^{\tau}\varphi(t)(J(t)+CE(t))^{\prime}\,\mathrm{d}t+C\int_{t_{0}}^{\tau}\varphi(t)(-E^{\prime}(t))^{\frac{2}{q+1}}\,\mathrm{d}t.$ (4.25)
For the first term on the right-hand side above, we perform integration by parts to calculate
$-\int_{t_{0}}^{\tau}\varphi(t)(J(t)+CE(t))^{\prime}\,\mathrm{d}t$ $=\Big{[}\varphi(t)(J(t)+CE(t))\Big{]}_{\tau}^{t_{0}}+\int_{t_{0}}^{\tau}\varphi^{\prime}(t)(J(t)+CE(t))\,\mathrm{d}t$
$\leq\varphi(t_{0})(J(t_{0})+CE(t_{0}))\leq C\varphi(0)E(0),$
by the monotonicity of $\varphi(t),E(t)$ and the same argument as before. For the second term on the right-hand side of (4.25), we take advantage of Young’s inequality for products to estimate
$C\int_{t_{0}}^{\tau}\varphi(t)(-E^{\prime}(t))^{\frac{2}{q+1}}\,\mathrm{d}t\leq\frac{1}{2}\int_{t_{0}}^{\tau}\varphi(t)^{\frac{q+1}{q-1}}\,\mathrm{d}t-C\int_{t_{0}}^{\tau}E^{\prime}(t)\,\mathrm{d}t\leq CE(0)+\frac{1}{2}\int_{t_{0}}^{\tau}\varphi(t)^{\frac{q+1}{q-1}}\,\mathrm{d}t.$
Especially, taking $\varphi(t)=E(t)^{\frac{q-1}{2}}$ gives $\varphi(t)E(t)=\varphi(t)^{\frac{q+1}{q-1}}=E(t)^{\frac{q+1}{2}}$. Thus we substitute the above two inequalities into (4.25) to conclude
$\int_{t_{0}}^{\tau}E(t)^{\frac{q+1}{2}}\,\mathrm{d}t\leq CE(0).$
The remaining proof is identical to that of Case 1.
5. Proof of Theorem 3.2
To establish Theorem 3.2, we require the following lemma.
Lemma 5.1.
Under assumptions (A4)–(A5), there exists a sufficiently small $\gamma\in(0,1)$ such that the term $(f\circ\nabla u)(t)$ defined by (2.5) satisfies
$\frac{1}{\gamma}G^{-1}\left(\frac{\gamma\,\mu(t)}{\xi(t)}\right),$ $q=1,$ (5.1)
$\frac{t}{\gamma}G^{-1}\left(\frac{\gamma\,\mu(t)}{t\,\xi(t)}\right),$ $q>1,$ (5.2)
where $\xi,G$ were introduced in assumption (A5) and
$\mu(t):=-(f^{\prime}\circ\nabla u)(t)\leq-2E^{\prime}(t).$ (5.3)
Proof.
Recall that the existence of the inverse function $G^{-1}$ was mentioned in Remark 3.1. Meanwhile, the inequality in (5.3) follows immediately from (2.9) and assumptions (A2)–(A4).
For $q=1$, let us fix any $t>0$ and set
$\psi(s):=\int_{\Omega}a|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x.$
Taking advantage of (2.8), (2.12) and (3.2) in Theorem 3.1, we invoke the non-negativity and the monotonicity of $E(t)$ to dominate
$\int_{0}^{t}\psi(s)\,\mathrm{d}s$ $\leq 2\|a\|_{\infty}\int_{0}^{t}\left\{\|\nabla u(t)\|_{2}^{2}+\|\nabla u(s)\|_{2}^{2}\right\}\mathrm{d}s\leq\frac{4\|a\|_{\infty}}{\ell}\int_{0}^{t}(\mathbb{E}(t)+\mathbb{E}(s))\,\mathrm{d}s$
$\leq C\int_{0}^{t}(E(t)+E(s))\,\mathrm{d}s\leq C\int_{0}^{t}E(s)\,\mathrm{d}s\leq C\int_{0}^{\infty}E(t)\,\mathrm{d}t\leq CE(0)<\infty.$
Then we can choose a sufficiently small $\gamma>0$ such that $\int_{0}^{t}\gamma\,\psi(s)\,\mathrm{d}s=1$, that is, $\gamma\,\psi(s)\,\mathrm{d}s$ is a probability measure on $(0,t)$. By assumption (A5) and especially the convexity of $G$, we utilize Jensen’s inequality to deduce
$\mu(t)$ $=\frac{1}{\gamma}\int_{0}^{t}(-f^{\prime}(t-s))\left(\gamma\int_{\Omega}a|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x\right)\mathrm{d}s\geq\frac{1}{\gamma}\int_{0}^{t}\,\xi(t-s)G(f(t-s))\,\gamma\,\psi(s)\,\mathrm{d}s$
$\geq\frac{\xi(t)}{\gamma}\int_{0}^{t}G(f(t-s))\,\gamma\,\psi(s)\,\mathrm{d}s\geq\frac{\xi(t)}{\gamma}G\left(\int_{0}^{t}f(t-s)\,\gamma\,\psi(s)\,\mathrm{d}s\right)=\frac{\xi(t)}{\gamma}G(\gamma(f\circ\nabla u)(t)).$
which indicates (5.1) by performing $G^{-1}$ on both sides above.
Next we deal with $q>1$. In the same manner, we obtain
$\frac{1}{t}\int_{0}^{t}\psi(s)\,\mathrm{d}s$ $\leq\frac{2\|a\|_{\infty}}{t}\int_{0}^{t}\left\{\|\nabla u(t)\|_{2}^{2}+\|\nabla u(s)\|_{2}^{2}\right\}\mathrm{d}s\leq\frac{4\|a\|_{\infty}}{\ell\,t}\int_{0}^{t}(\mathbb{E}(t)+\mathbb{E}(s))\,\mathrm{d}s$
$\leq\frac{C}{t}\int_{0}^{t}E(s)\,\mathrm{d}s\leq\frac{C}{t}\int_{0}^{t}E(0)\,\mathrm{d}s\leq CE(0)<\infty.$
Then again we can choose a sufficiently small $\gamma>0$ such that $\frac{\gamma}{t}\int_{0}^{t}\psi(s)\,\mathrm{d}s=1$. Parallel to the above argument, it follows that
$\mu(t)$ $=\frac{t}{\gamma}\int_{0}^{t}(-f^{\prime}(t-s))\left(\frac{\gamma}{t}\int_{\Omega}a|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x\right)\mathrm{d}s$
$\geq\frac{t}{\gamma}\int_{0}^{t}\,\xi(t-s)G(f(t-s))\,\frac{\gamma}{t}\,\psi(s)\,\mathrm{d}s\geq\frac{\xi(t)t}{\gamma}\int_{0}^{t}G(f(t-s))\,\frac{\gamma}{t}\,\psi(s)\,\mathrm{d}s$
$\geq\frac{\xi(t)t}{\gamma}G\left(\frac{\gamma}{t}\,\int_{0}^{t}f(t-s)\,\psi(s)\,\mathrm{d}s\right)=\frac{\xi(t)t}{\gamma}G\left(\frac{\gamma}{t}(f\circ\nabla u)(t)\right),$
which indicates (5.2). ∎
Now we are in a position to prove Theorem 3.2. As before, we divide the cases of $q=1$ and $q>1$ separately.
Case 1. In the case of $q=1$, first we use Hölder’s inequality to deal with
$\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x$ $\leq\int_{\Omega}a\int_{0}^{t}f(s)\,\mathrm{d}s\int_{0}^{t}f(t-s)|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x\mathrm{d}s$
$\leq\frac{\lambda_{0}-\ell}{\|a\|_{\infty}}(f\circ\nabla u)(t).$ (5.4)
Combining (4.19), (4.21), (4.23) with (5.4), we derive the necessary estimate as follows
$L^{\prime}(t)\leq-\epsilon\,E(t)+C(f\circ\nabla u)(t)-CE^{\prime}(t).$
Using the estimate (5.1) in Lemma 5.1 and putting $L_{1}(t):=L(t)+CE(t)\sim E(t)$, we get
$L_{1}^{\prime}(t)\leq-\epsilon\,E(t)+CG^{-1}\left(\frac{\gamma\,\mu(t)}{\xi(t)}\right).$
Now we pick a constant $\varepsilon_{1}\in(0,f(0))$ and define
$S(t):=G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)L_{1}(t),$
which satisfies $S(t)\sim E(t)$. Noting that $G^{\prime\prime}(t)\geq 0$ and $E^{\prime}(t)\leq 0$, we estimate
$S^{\prime}(t)$ $=G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)L_{1}^{\prime}(t)+\varepsilon_{1}\frac{E^{\prime}(t)}{E(0)}G^{\prime\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)L_{1}(t)$
$\leq-\epsilon\,E(t)G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)+CG^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)G^{-1}\left(\frac{\gamma\,\mu(t)}{\xi(t)}\right).$ (5.5)
Next, let $G^{*}$ be the convex conjugate of $G$ in the sense of Young (see [2]), that is,
$G^{*}(s):=s\,(G^{\prime})^{-1}(s)-G((G^{\prime})^{-1}(s)),\quad s\in[0,G^{\prime}(f(0))].$
On the other hand, it is known that $G^{*}$ satisfies the generalized Young’s inequality
$AB\leq G^{*}(A)+G(B),\quad A\in[0,G^{\prime}(f(0))],\ B\in[0,f(0)].$ (5.6)
Taking $A=G^{\prime}(\varepsilon_{1}\frac{E(t)}{E(0)})$ and $B=G^{-1}(\frac{\gamma\,\mu(t)}{\xi(t)})$ in (5.6), we can estimate (5.5) as
$S^{\prime}(t)$ $\leq-\epsilon\,E(t)G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)+CG^{*}\left(G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)\right)+C\frac{\gamma\,\mu(t)}{\xi(t)}$
$\leq-\epsilon\,E(t)G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)+C\varepsilon_{1}\frac{E(t)}{E(0)}G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)+C\frac{\gamma\,\mu(t)}{\xi(t)}.$
Multiplying both sides of the above estimate by $\xi(t)$, we use (5.3) and the facts that $\varepsilon_{1}\frac{E(t)}{E(0)}<f(0)$ to derive
$\xi(t)S^{\prime}(t)$ $\leq-\epsilon\,\xi(t)E(t)G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)+C\xi(t)\varepsilon_{1}\frac{E(t)}{E(0)}G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)+C\gamma\,\mu(t)$
$\leq-(\epsilon\,E(0)-C\varepsilon_{1})\xi(t)\frac{E(t)}{E(0)}G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)-CE^{\prime}(t).$
Further setting $S_{1}(t):=\xi(t)S(t)+CE(t)$, we know $S_{1}(t)\sim E(t)$, that is, there exist constants $\alpha_{1},\alpha_{2}>0$ such that
$\alpha_{1}S_{1}(t)\leq E(t)\leq\alpha_{2}S_{1}(t).$ (5.7)
With a suitable choice of $\varepsilon_{1}>0$, we have
$S_{1}^{\prime}(t)\leq-\alpha_{3}\xi(t)\frac{E(t)}{E(0)}G^{\prime}\left(\varepsilon_{1}\frac{E(t)}{E(0)}\right)=-\alpha_{3}\xi(t)G_{2}\left(\frac{E(t)}{E(0)}\right),$ (5.8)
where $\alpha_{3}:=\epsilon\,E(0)-C\varepsilon_{1}>0$ and $G_{2}(t)=t\,G^{\prime}(\varepsilon_{1}t)$. By
$G_{2}^{\prime}(t)=G^{\prime}(\varepsilon_{1}t)+\varepsilon_{1}t\,G^{\prime\prime}(\varepsilon_{1}t)$
and the properties of $G$, we see that $G_{2},G_{2}^{\prime}>0$ on $[0,f(0)]$. Finally introducing $S_{2}(t):=\frac{\alpha_{1}S_{1}(t)}{E(0)}\sim E(t)$, we employ (5.7)–(5.8) to obtain
$S_{2}^{\prime}(t)\leq-k_{1}{\xi(t)}G_{2}\left(\frac{E(t)}{E(0)}\right)\leq-k_{1}\xi(t)G_{2}\left(\frac{\alpha_{1}S_{1}(t)}{E(0)}\right)=-k_{1}\xi(t)G_{2}(S_{2}(t)),$
where $k_{1}:=\frac{\alpha_{1}\alpha_{3}}{E(0)}$. Therefore, we divide both sides of the above inequality by $G_{2}(S_{2}(t))$ and integrate over some interval $(t_{0},t)\subset\mathbb{R}_{+}$ to deduce
$k_{1}\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s\leq-\int_{t_{0}}^{t}\frac{S^{\prime}_{2}(s)}{G_{2}(S_{2}(s))}\,\mathrm{d}s=\int_{\varepsilon_{1}S_{2}(t)}^{\varepsilon_{1}S_{2}(t_{0})}\frac{1}{s\,G^{\prime}(s)}\,\mathrm{d}s$
by some appropriate change of variables. Fixing $t_{0}$ such that $\varepsilon_{1}S_{2}(t_{0})=f(0)$ and recalling the definition of $G_{1}(t)$ in Theorem 3.2, we can rewrite the above inequality as
$G_{1}(\varepsilon_{1}S_{2}(t))\geq k_{1}\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s.$
Owing to the fact that $G_{1}$ is strictly decreasing on $[0,f(0)]$, we conclude
$S_{2}(t)\leq\frac{1}{\varepsilon_{1}}G_{1}^{-1}\left(k_{1}\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s\right),$
which eventually implies (3.4) by the equivalence of $S_{2}(t)$ and $E(t)$.
Case 2. Now let us treat the case of $q>1$. First we combine the estimate (5.2) in Lemma 5.1 with (4.24) and (5.4) to obtain
$L^{\prime}(t)\leq-\epsilon\,E(t)+C\frac{t}{\gamma}G^{-1}\left(\frac{\gamma\,\mu(t)}{t\,\xi(t)}\right)-CE^{\prime}(t)+C(-E^{\prime}(t))^{\frac{2}{q+1}}.$
In an analogous manner as before, we multiply both sides of the above inequality by $E(t)^{\frac{q-1}{2}}$ and employ Young’s inequality for products to estimate
$E(t)^{\frac{q-1}{2}}L^{\prime}(t)$ $\leq-\epsilon\,E(t)^{\frac{q+1}{2}}+CE(0)^{\frac{q-1}{2}}t\,G^{-1}\left(\frac{\gamma\,\mu(t)}{t\,\xi(t)}\right)-CE^{\prime}(t)E(t)^{\frac{q-1}{2}}+C(-E^{\prime}(t))^{\frac{2}{q+1}}E(t)^{\frac{q-1}{2}}$
$\leq-\frac{\epsilon}{2}E(t)^{\frac{q+1}{2}}+C\,t\,G^{-1}\left(\frac{\gamma\,\mu(t)}{t\,\xi(t)}\right)-CE^{\prime}(t)E(t)^{\frac{q-1}{2}}-CE^{\prime}(t).$
Introducing an auxiliary function
$R(t):=E(t)^{\frac{q-1}{2}}L(t)+CE(t)^{\frac{q+1}{2}}+CE(t),$
we have $R(t)\sim E(t)$ and
$R^{\prime}(t)\leq-\frac{\epsilon}{2}E(t)^{\frac{q+1}{2}}+C\,t\,G^{-1}\left(\frac{\gamma\,\mu(t)}{t\,\xi(t)}\right).$ (5.9)
Next, we pick a constant $\varepsilon_{1}\in(0,f(0))$ and define
$R_{1}(t):=G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)R(t),$
which satisfies $R_{1}(t)\sim E(t)$. Noticing $G^{\prime\prime}(t)\geq 0$ and $E^{\prime}(t)\leq 0$, we use (5.9) to deduce
$R_{1}^{\prime}(t)$ $\leq G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)R^{\prime}(t)$
$\leq-\frac{\epsilon}{2}E(t)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)+C\,t\,G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)G^{-1}\left(\frac{\gamma\,\mu(t)}{t\,\xi(t)}\right).$
Here we apply (5.6) with
$A=G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right),\quad B=G^{-1}\left(\frac{\gamma\,\mu(t)}{t\,\xi(t)}\right)$
to the above inequality to further bound
$R_{1}^{\prime}(t)$ $\leq-\frac{\epsilon}{2}E(t)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)+C\,t\,G^{*}\left(G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)\right)+C\frac{\gamma\,\mu(t)}{\xi(t)}$
$\leq-\frac{\epsilon}{2}E(t)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)+C\varepsilon_{1}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)+C\frac{\gamma\,\mu(t)}{\xi(t)}.$
Multiplying both sides of the above estimate by $\xi(t)$, we use (5.3) and the facts that
$\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}<f(0)$
to derive
$\xi(t)R_{1}^{\prime}(t)$ $\leq-\frac{\epsilon}{2}\xi(t)E(t)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)$
$\quad\,+C\xi(t)\varepsilon_{1}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)+C\gamma\,\mu(t)$
$\leq-\varepsilon_{2}\xi(t)\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)-CE^{\prime}(t),$ (5.10)
where
$\varepsilon_{2}:=\frac{\epsilon}{2}E(0)^{\frac{q+1}{2}}-C\varepsilon_{1}$
and we can choose $\varepsilon_{1}>0$ sufficiently small such that $\varepsilon_{2}>0$.
By further putting $R_{2}(t):=\xi(t)R_{1}(t)+CE(t)$ as $S_{1}(t)$ in Case 1, again we have $R_{2}(t)\sim E(t)$. Then we employ $\xi^{\prime}(t)\leq 0$ and (5.10) to estimate $R_{2}^{\prime}(t)$ as
$R_{2}^{\prime}(t)=\xi^{\prime}(t)R_{1}(t)+\xi(t)R_{1}^{\prime}(t)+CE^{\prime}(t)\leq-\varepsilon_{2}\xi(t)\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)\leq 0.$
Then obviously $R_{2}^{\prime}(t)\leq 0$ and integrating the above inequality over some interval $(t_{0},t)\subset\mathbb{R}_{+}$ yields
$\varepsilon_{2}\int_{t_{0}}^{t}\xi(s)\left(\frac{E(s)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{s}\left(\frac{E(s)}{E(0)}\right)^{\frac{q+1}{2}}\right)\mathrm{d}s\leq-\int_{t_{0}}^{t}R_{2}^{\prime}(s)\,\mathrm{d}s\leq R_{2}(t_{0}).$
By the monotonicity of $G^{\prime}(t),G^{\prime\prime}(t)$ and $E(t)$, the function
$t\longmapsto\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)$
turns out to be non-increasing. Therefore, we arrive at
$\varepsilon_{2}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s\leq R_{2}(t_{0}).$
Dividing both sides of the above estimate by $t$ and recalling the definition of $G_{2}(t)$ in Theorem 3.2, we conclude
$\varepsilon_{2}G_{2}\left(\frac{1}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s\leq\frac{R_{2}(t_{0})}{t}.$
Consequently, owing to the fact that $G_{2}$ is strictly increasing in $\overline{\mathbb{R}_{+}}$, we obtain the desired estimate (3.5) by putting $k_{2}:=R_{2}(t_{0})/\varepsilon_{2}$. The proof of Theorem 3.2 is completed.
Acknowledgements
The first author is supported by China Scholarship Council. The second author is supported by JSPS KAKENHI Grant Numbers JP22K13954 and JP23KK0049.

### v2

@@
 The main purpose of this manuscript is to investigate the energy decay rate of the following initial-boundary value problem for a nonlocal viscoelastic wave equation with variable coefficients, a nonlinear damping term and a polynomial nonlinear term:
-{$&u_{tt}-\mathrm{div}(A(x)\nabla u)+\int_{0}^{t}f(t-s)\,\mathrm{div}(a(x)\nabla u(s))\,\mathrm{d}s\\ &\quad+b(x)h(u_{t})=k(x)|u|^{p-2}u\end{aligned}&\quad&\mbox{in }\Omega\times\mathbb{R}_{+},\\ &u=u^{0},\ u_{t}=u^{1}&\quad&\mbox{in }\Omega\times\{0\},\\ &u=0&\quad&\mbox{on }\partial\Omega\times\mathbb{R}_{+},\end{aligned}\right.$ (1.1)
+$\left\{\begin{aligned} &\begin{aligned} &u_{tt}-\mathrm{div}(A(x)\nabla u)+\int_{0}^{t}f(t-s)\,\mathrm{div}(a(x)\nabla u(s))\,\mathrm{d}s\\ &\quad+b(x)h(u_{t})=k(x)|u|^{p-2}u\end{aligned}&\quad&\mbox{in }\Omega\times\mathbb{R}_{+},\\ &u=u^{0},\ u_{t}=u^{1}&\quad&\mbox{in }\Omega\times\{0\},\\ &u=0&\quad&\mbox{on }\partial\Omega\times\mathbb{R}_{+},\end{aligned}\right$.
+(1.1)
 where the initial displacement $u^{0}\in H_{0}^{1}(\Omega)$ and the initial velocity $u^{1}\in L^{2}(\Omega)$.
@@
 Moreover, the exponent constant $p$ satisfies
-[…]
+$\left\{\begin{aligned} &p>2,&\quad&n=1,2,\\ &2<p\leq\frac{2n-2}{n-2},&\quad&n\geq 3.\end{aligned}\right$.
 Detailed assumptions on $A(x),f(t),a(x),b(x)$ will be specified in Section 2.
@@
 As a pioneering work, Webb [19] first investigated the well-posedness of solutions through the operator theory and examined the long-term behavior of these solutions using Lyapunov stability techniques for the semilinear wave equation with strong damping and a forcing term.
-Gazzola and Squassina [8] studied the semilinear wave equation with strong frictional damping and established both existence and non-existence results for solutions when the initial energy does not exceed the depth of the potential well.
+Gazzola and Squassina [8] studied the semilinear wave equation with strong frictional damping and established both existence and nonexistence results for solutions when the initial energy does not exceed the depth of the potential well.
 Further, they discovered the blow-up phenomena by potential well methods along with Levine’s concavity techniques.
-Subsequently, Ma and Fang [15] demonstrated both existence and non-existence of global weak solutions and derived energy decay estimates by combining potential well methods with a logarithmic Sobolev inequality applicable to the viscoelastic wave equations.
+Subsequently, Ma and Fang [15] demonstrated both existence and nonexistence of global weak solutions and derived energy decay estimates by combining potential well methods with a logarithmic Sobolev inequality applicable to the viscoelastic wave equations.
 Lian and Xu [13] further explored global existence, energy decay and infinite-time blow-up using methodologies similarly to [15].
@@
 The remainder of this paper is organized as follows.
-In Section 2, we specify assumptions on coefficients and prepare prelimineary lemmas including some elementary estimates of energies the global well-posedness of solutions.
+In Section 2, we specify assumptions on coefficients and prepare preliminary lemmas including some elementary estimates of energies the global well-posedness of solutions.
 In Section 3, we state the two main results of this article, namely, the polynomial energy decay in Theorem 3.1 as well as a refined decay estimate in Theorem 3.2 under a weak convexity assumption on the kernel function $f(t)$.
@@
 In this section, we fix notations and introduce some basic definitions, important lemmas and function spaces for the sake of the statements and proofs of our main results.
-[…]
+Throughout this article, we denote the norm of the Lebesgue space $L^{p}(\Omega)$ ($1\leq p\leq\infty$) by $\|\cdot\|_{p}$, and the inner product of $L^{2}(\Omega)$ by (,)($\cdot\,$$\cdot\,$).
 2.1. Some assumptions and definitions
@@
 $a+b\geq\kappa>0\mbox{ in }\Omega.$
-(A3) The function $h\in C(\mathbb{R})$ is non-decreasing and satisfies
+(A3) The function $h\in C(\mathbb{R})$ is nondecreasing and satisfies
 $h(s)s\geq 0,\quad\forall\,s\in\mathbb{R}.$
@@
 $\begin{cases}c_{1}|s|\leq|h(s)|\leq c_{2}|s|,&|s|>1,\\ c_{3}|s|^{q}\leq|h(s)|\leq c_{4}|s|^{1/q},&|s|\leq 1.\end{cases}$
-(A4) The function $f\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is non-increasing and satisfies
+(A4) The function $f\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is nonincreasing and satisfies
 $f(0)>0,\quad\lambda_{0}-\|a\|_{\infty}\int_{0}^{\infty}f(t)\,\mathrm{d}t:=\ell>0.$ (2.2)
@@
 $E^{\prime}(t)=\frac{1}{2}(f^{\prime}\circ\nabla u)(t)-\frac{f(t)}{2}\int_{\Omega}a|\nabla u(t)|^{2}\,\mathrm{d}x-\int_{\Omega}b\,h(u_{t}(t))u_{t}(t)\,\mathrm{d}x.$ (2.9)
-Thanks to assumptions (A2)–(A4), we see that $E^{\prime}(t)\leq 0$, that is, the energy $E(t)$ is non-increasing.
-2.2. Some lemmas and well-posedness
+Thanks to assumptions (A2)–(A4), we see that $E^{\prime}(t)\leq 0$, that is, the energy $E(t)$ is nonincreasing.
+$2.$ Some lemmas and well-posedness
 Lemma 2.1 (see [1]).
@@
 $\|w\|_{r}^{r}\leq B_{r}\|\nabla w\|_{2}^{r},\quad\forall\,w\in H_{0}^{1}(\Omega).$
-Lemma 2.2.
+Lemma $2.$
 Under assumptions (A1)–(A4), define
@@
 $f^{\prime}(t)\leq-\xi(t)G(f(t)),\quad t>0.$
-Here $\xi\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is non-increasing and $G\in C^{2}([0,f(0)];\overline{\mathbb{R}_{+}})$ is a strictly increasing and strictly convex function satisfying $G(0)=G^{\prime}(0)=0$.
+Here $\xi\in C^{1}(\overline{\mathbb{R}_{+}};\mathbb{R}_{+})$ is nonincreasing and $G\in C^{2}([0,f(0)];\overline{\mathbb{R}_{+}})$ is a strictly increasing and strictly convex function satisfying $G(0)=G^{\prime}(0)=0$.
 Remark 3.1 (see [16]).
@@
 According to the assumptions on $G$ in (A5) and the above definitions of $G_{1},G_{2}$, we see that $G_{1}$ is strictly decreasing and defines a bijection from $[0,f(0)]$ to $\overline{\mathbb{R}_{+}}$, and $G_{2}$ is strictly increasing and defines a bijection from $\overline{\mathbb{R}_{+}}$ to $\overline{\mathbb{R}_{+}}$.
-Remark 3.3.
+Remark $3.$
 Theorems 3.13.2 still hold for $p=2$, i.e., the polynomial nonlinearity in (1.1) degenerates to the linear one. The proofs are rather analogous and thus we omit them here.
@@
 Example 3.2.
-Let $q=1$ and $f(t)=\alpha\exp(-t^{\beta})$, where $\beta\in(0,1)$ and $a>0$ are again small constants so that $f$ satisfies (A4). Then this time $f$ satisfies (A5) with $\xi(t)=1$ and
+Let $q=1$ and $f(t)=\alpha\exp(-t^{\beta})$, where $\beta\in(0,1)$ and $\alpha>0$ are again small constants so that $f$ satisfies (A4). Then this time $f$ satisfies (A5) with $\xi(t)=1$ and
 $G(t)=\frac{\beta\,t}{(\ln(\alpha/t))^{1/\beta-1}}.$
@@
 $G^{\prime}(t)=\frac{(1-\beta)+\beta\ln(\alpha/t)}{(\ln(\alpha/t))^{1/\beta}},\quad G^{\prime\prime}(t)=\frac{(1-\beta)(\ln(\alpha/t)+1/\beta)}{(\ln(\alpha/t))^{\frac{1}{\beta+1}}},$
-we see that $G$ also achieves the requirement in (A5) on $[0,\alpha]$. Then we deal with $G_{1}$ as
+we see that $G$ also achieves the requirement in (A5) on $0,\alpha]$. Then we deal with $G_{1}$ as
 $G_{1}(t)$ $=\int_{t}^{\alpha}\frac{1}{s\,G^{\prime}(s)}\,\mathrm{d}s=\int_{t}^{\alpha}\frac{(\ln(\alpha/s))^{1/\beta}}{s(1-\beta+\beta\ln(\alpha/s))}\,\mathrm{d}s=\int_{0}^{\ln(\alpha/t)}\frac{\tau^{1/\beta}}{1-\beta+\beta\tau}\,\mathrm{d}\tau$
@@
 Then it follows from (3.4) that $E(t)\leq c\exp(-c\,t^{\beta})$ with some constant $c>0$.
-Example 3.3.
+Example $3.$
 Now let us consider $q>1$ and $f(t)=\alpha(1+t)^{-\beta}$, where $\alpha>0$ and $\beta>1$ are suitably chosen so that (A4) holds.
@@
 $I_{1}(t):=\int_{\Omega}a\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla u(s)\,\mathrm{d}s\mathrm{d}x,\quad I_{2}(t):=-\int_{\Omega}b\,h(u_{t}(t))u(t)\,\mathrm{d}x.$
-For $I_{1}(t)$, we separate $u(s)=u(t)+(u(s)-u(t))$ and use assumptions (A2) and (A4) to estimate
+For $I_{1}(t)$, we separate u(s)=u(t)+(u(s)u(t))u(s)=u(t)+(u(s)-u(t)) and use assumptions (A2) and (A4) to estimate
 $I_{1}(t)$ $=\int_{\Omega}a\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla u(t)\,\mathrm{d}s\mathrm{d}x+\int_{\Omega}a\nabla u(t)\cdot\int_{0}^{t}f(t-s)\nabla(u(s)-u(t))\,\mathrm{d}s\mathrm{d}x$
@@
 $=\int_{\Omega}\left\{-\mathrm{div}(A\nabla u(t))+\int_{0}^{t}f(t-s)\,\mathrm{div}(a\nabla u(s))\,\mathrm{d}s+b\,h(u_{t}(t))-k\,|u(t)|^{p-2}u(t)\right\}$
-×$\qquad\>\>\times\widetilde{a}\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x+I_{9}(t)+I_{10}(t)$
+$\qquad\>\>\times\widetilde{a}\int_{0}^{t}f(t-s)(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x+I_{9}(t)+I_{10}(t)$
 $=\sum_{i=3}^{10}I_{i}(t),$
@@
 $\leq\frac{\varepsilon}{4}\|\nabla u(t)\|_{2}^{2}+\frac{C\|\nabla a\|_{\infty}^{2}\widetilde{\lambda}}{\varepsilon}\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x.$
-For $I_{5}(t)$, we separate $u(s)=u(t)+(u(s)-u(t))$ and use assumptions (A2) and (A4) to estimate
+For $I_{5}(t)$, we separate u(s)=u(t)+(u(s)u(t))u(s)=u(t)+(u(s)-u(t)) and use assumptions (A2) and (A4) to estimate
 $|I_{5}(t)|$ $=\left|\int_{\Omega}\widetilde{a}\int_{0}^{t}f(t-s)a\nabla(u(s)-u(t)+u(t))\,\mathrm{d}s\cdot\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\mathrm{d}x\right|$
@@
 $\leq\varepsilon B_{2(p-1)}\eta_{1}\|\nabla u(t)\|_{2}^{2}+C(\varepsilon)\int_{\Omega}a\left|\int_{0}^{t}f(t-s)\nabla(u(t)-u(s))\,\mathrm{d}s\right|^{2}\mathrm{d}x,$ (4.18)
-where we used Lemma 2.1 with $r=2(p-1)$ and recall $K=\|k\|_{\infty}$. For the first term in (4.18), we combine (2.12) with (2.8) to deduce
+where we used Lemma 2.1 with r=2(p1)r=2(p-1) and recall $K=\|k\|_{\infty}$. For the first term in (4.18), we combine (2.12) with (2.8) to deduce
 $\|\nabla u(t)\|_{2}^{2(p-2)}\leq\left(\frac{2\,\mathbb{E}(t)}{\ell}\right)^{p-2}\leq\left(\frac{2(1+\widetilde{C})E(0)}{\ell}\right)^{p-2}=\eta_{1}.$
@@
 $J^{\prime}(t)+\epsilon\,E(t)\leq C\left\{-E^{\prime}(t)+(-E^{\prime}(t))^{\frac{2}{q+1}}\right\}.$ (4.24)
-Now introduce an auxiliary function $\varphi:\overline{\mathbb{R}_{+}}\longrightarrow\overline{\mathbb{R}_{+}}$ which is non-increasing and locally absolutely continuous.
+Now introduce an auxiliary function $\varphi:\overline{\mathbb{R}_{+}}\longrightarrow\overline{\mathbb{R}_{+}}$ which is nonincreasing and locally absolutely continuous.
 For the constant $t_{0}>0$ in Case 1 and any $\tau>t_{0}$, we multiply (4.24) by $\varphi(t)$ and integrate over $(t_{0},\tau)$ to deduce
@@
 $\psi(s):=\int_{\Omega}a|\nabla(u(t)-u(s))|^{2}\,\mathrm{d}x.$
-Taking advantage of (2.8), (2.12) and (3.2) in Theorem 3.1, we invoke the non-negativity and the monotonicity of $E(t)$ to dominate
+Taking advantage of (2.8), (2.12) and (3.2) in Theorem 3.1, we invoke the nonnegativity and the monotonicity of $E(t)$ to dominate
 $\int_{0}^{t}\psi(s)\,\mathrm{d}s$ $\leq 2\|a\|_{\infty}\int_{0}^{t}\left\{\|\nabla u(t)\|_{2}^{2}+\|\nabla u(s)\|_{2}^{2}\right\}\mathrm{d}s\leq\frac{4\|a\|_{\infty}}{\ell}\int_{0}^{t}(\mathbb{E}(t)+\mathbb{E}(s))\,\mathrm{d}s$
@@
 $t\longmapsto\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)$
-turns out to be non-increasing. Therefore, we arrive at
+turns out to be nonincreasing. Therefore, we arrive at
 $\varepsilon_{2}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}G^{\prime}\left(\frac{\varepsilon_{1}}{t}\left(\frac{E(t)}{E(0)}\right)^{\frac{q+1}{2}}\right)\int_{t_{0}}^{t}\,\xi(s)\,\mathrm{d}s\leq R_{2}(t_{0}).$